OpenStudy (anonymous):

triangle PQR is right angled at P. The points P(2;6), Q(-1;2) and R(6;k) Determine K

OpenStudy (amistre64):

helps to draw a general setup to get the juices flowing

OpenStudy (amistre64):



OpenStudy (amistre64):

id "move" P to the origin (2-2,6-6) = (0,0) and the others move in kind Q=(-1-2,2-6) = (-3,-4) R=(6-2,k-6) = (4,k')

OpenStudy (amistre64):

can you find the slope from the P to Q now?

OpenStudy (anonymous):

Yeah i have worked out the gradient it's 4 and since it's perpendicular the gradient of PR is -1/4 so thats where i'm stuck

OpenStudy (amistre64):

another method if youknow what a dot product is might be simpler ....